OpenStudy (anonymous):

ln [(3+4x^2)^1/2] find the domain and derivative

OpenStudy (anonymous):

\[\large y = \ln (3+4x^2)^\frac {1}{2}\] \[\large y' = \frac {[(3+4x^2)^\frac {1}{2}]'}{(3+4x^2)^\frac {1}{2}}\] \[\large y' = \frac {\frac {1}{2}(3+4x^2)^{-1/2}(8x)}{(3+4x^2)^\frac {1}{2}}\] simplify the rest...

OpenStudy (anonymous):

I see where I messed up, thanks!

OpenStudy (anonymous):

yw... btw, do you know the domain?

OpenStudy (anonymous):

negative to positive infinity right?

OpenStudy (anonymous):

correct... all reals.. :) good work.